OpenStudy (anonymous):

Let vector AB=(6,-2,3)and vector AC= (-2,-3,2). How do I find vector BC?

OpenStudy (raden):

AB = b-a AC = c-a BC = c-b = (c-a) - (b-a) this is a key to slove this problem

OpenStudy (raden):

* solve

OpenStudy (anonymous):

so, does BC equal (-8,-1,-1)?

OpenStudy (raden):

yes, u are right sorry, too late replay this. my connection very slow :(

OpenStudy (anonymous):

oh it's alright. thankyou